# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v3

@@ 3.2.2. From GAN- to SAN-based discriminator
We apply the slicing adversarial network (SAN) framework (Takida et al., 2024) to further enhance the motion VAE, based on a prior report showing SAN-based models perform better than the GAN counterparts. The impact of this replacement on motion generation is discussed in Section 4, while a detailed explanation of SAN is provided in the Appendix.
3.2.3. Architectures
@@
We use a standard CNN-based architecture in the motion VAE encoder $q_{\bm{\eta}}$, decoder $g_{\bm{\psi}}$ and discriminator $f_{\bm{\psi}}$, consisting of 1D convolution, a residual block, and Leaky ReLU. For temporal downsampling and upsampling, we use stride 2 convolution and nearest interpolation, respectively. Specifically, the motion sequence $\bm{x}\in\mathbb{R}^{N\times L}$ is encoded into a latent vector $\bm{z}\in\mathbb{R}^{d_{z}\times d_{l}}$ with downsampling ratio of $d_{l}=L/4$. This architecture is inspired by (Esser et al., 2021; Zhang et al., 2023b).
3.3. Stage 2: Motion Latent Diffusion
3.3.1. Text-conditional motion generation
@@
In this section, we train a text-conditioned diffusion model on the low-dimensional motion latent space obtained by the autoencoder learned in the stage 1 (Section 3.2). Using the trained model, we perform motion generation conditioned on text. First, we define a time-dependent sequence $\bm{z}_{0},\bm{z}_{1},\dots,\bm{z}_{t},\dots,\bm{z}_{T}\in\mathbb{R}^{d_{z}\times d_{l}}$ (starting from the VAE encoder output $\bm{z}_{0}=\bm{z}\sim q_{\bm{\eta}}(\bm{z}|\bm{x})$), which is derived from the following Markov diffusion process in the latent space:
(8) $q(\bm{z}_{t}|\bm{z}_{t-1})=\mathcal{N}(\sqrt{\alpha}_{t}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$
where $T>0$ and the constant $\alpha_{t}\in(0,1)$ is a pre-defined noise-scheduling parameter that determines the forward process. The forward process allows for the sampling of $\bm{z}_{t}$ at an arbitrary time step $t$ in a closed form: $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}\coloneq\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$.
@@
As our goal is text-to-motion generation, our interest is in the conditional distribution $p(\bm{z}|\bm{c})$ given the text prompt $\bm{c}$. Here, similar to many text-conditioned latent diffusion models (Rombach et al., 2022; Chen et al., 2023), we train the conditional model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ conditioned on the output of a text encoder $\tau(\bm{c})$, using the following objective function:
@@
(9) $\mathcal{J}_{\text{cLDM}}(\theta)=\mathbb{E}_{\{\bm{z}_{0},\bm{c}\},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))\|_{2}^{2}\right],$
@@
where $\bm{z}_{0}$ and $\bm{c}$ are drawn from the joint empirical distribution. In addition, as done in prior works, we adopt classifier-free guidance (Ho and Salimans, 2022) and train the model unconditionally, i.e., without a text prompt, with a certain probability during training.
@@
During inference (text-to-motion generation), the trained diffusion model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ is used to generate $\bm{z}_{0}$ through a denoising process conditioned on the text prompt $\bm{c}$. We adopt the sampling scheme of DDIM (Song et al., 2021a) with trailing sample steps (Lin et al., 2024), in which each sampling step is defined as:
@@
(10) $\ +\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))+\sigma_{t}\epsilon,$
@@
where $\sigma_{t}>0$ determines the stochasticity of the sampling process, and the sampling process becomes deterministic when $\sigma_{t}=0$. The part $(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c})))/\sqrt{\bar{\alpha}_{t}}$ in the first term corresponds to a direct estimate of the clean latent $\bm{z}_{0}$ from the noisy sample $\bm{z}_{t}$ using the diffusion model based on Tweedie’s formula (Efron, 2011); this estimate is denoted as $\bm{z}_{0|t}$.
In actual inference, we use the estimated $z_{0}$ and the VAE decoder trained in stage 1 to obtain $g_{\bm{\psi}}(\bm{z}_{0})$ as the generated motion sequences. Variable length motion generation is then achieved by clipping a part of $g_{\bm{\psi}}(\bm{z}_{0})$ where the activation variable $a^{i}$ introduced in Section 3.1 satisfies $a^{i}<\delta\ (0<\delta<1)$. The effect of this approach is discussed in Section 4.1.
3.3.2. Architecture
@@
We employ a diffusion transformer (DiT)-based architecture in our stage 2 model. The transformer used follows a standard structure of stacked blocks consisting of an attention layer and gated multilayer perceptrons (MLP) connected in series, with skip connections around each. Additionally, layer normalization is employed on the inputs of both the attention layer and the MLP. At the input and output of the transformer, linear mapping is used to convert from the latent dimension of the stage 1 model to the embedded dimension of the transformer. Text CLIP embedding is also added as input to the transformer, along with an embedding describing the current time step of the diffusion process. This architecture is inspired by (Evans et al., 2024), which established SOTA performance in audio generation.
3.4. Controllable Motion Generation on Latent Diffusion Sampling
In this section, we present a guided generation framework that leverages the pre-trained motion latent diffusion model (Section 3.3) for conditional motion generation and editing tasks without extra training. Major training-free methods such as DPS (Chung et al., 2023), FreeDoM (Yu et al., 2023), and MPGD (He et al., 2024) are based on the fact that the conditional score function can be decomposed into two additive terms: the unconditional score function and the log-likelihood term. Specifically, for a new condition $\bm{y}$, we have
(11) $\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t}|\bm{c},\bm{y})=\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t}|\bm{c})+\nabla_{\bm{z}_{t}}\log p(\bm{y}|\bm{z}_{t},\bm{c}),$
@@ 3.4. Controllable Motion Generation on Latent Diffusion Sampling
which is derived from Bayes’ rule. The conditional generation based on the aforementioned property can be regarded as a sequential procedure implemented as follows. First, a denoised sample $\bm{z}_{t-1}$ is obtained from $\bm{z}_{t}$ by the sampling step in Equation (10), without considering the given condition $\bm{y}$ (the first term in Equation (10)). Subsequently, $\bm{z}_{t-1}$ is further updated using the gradient of the log-likelihood term with respect to $\bm{z}_{t}$ (the second term).
@@
The challenge here is that the log-likelihood term is computed based on noisy samples $\bm{z}_{t}$. In the classifier-guided diffusion (Song et al., 2021b), time-dependent classifiers for $\bm{z}_{t}$ have to be trained, which requires additional training. In contrast, in training-free methods such as DPS, FreeDoM, and MPGD, this term is approximated with the current clean data estimate $\bm{z}_{0|t}$ and a loss function $\mathfrak{L}(\bm{x};\bm{y})$ defined for clean data. This loss function can be flexibly set depending on the task. For example, in inverse problems of the form $\bm{y}=\mathcal{A}(\bm{x})$, where $\mathcal{A}$ is a differentiable function with respect to $\bm{x}$, the loss function can be set as $\|\bm{y}-\mathcal{A}(\bm{x}_{0|t})\|_{2}^{2}$, where $\bm{x}_{0|t}=g_{\bm{\psi}}(\bm{z}_{0|t})$ is a clean data estimate in the original data domain and obtained through the VAE decoder. In FreeDoM’s FaceID-guided generation, the $\ell_{2}$ distance of features obtained by inputting facial images into a facial recognition model is used as the loss.
Here, we adopt MPGD (He et al., 2024), a fast yet high-quality guidance method applicable to latent diffusion models. Following the denoising step, it updates the denoised sample based on the loss function $\mathfrak{L}(\bm{x};\bm{y})$ as follows:
(12) $\bm{z}_{t-1}\leftarrow\bm{z}_{t-1}-\rho_{t}\sqrt{\bar{\alpha}_{t-1}}\nabla_{\bm{z}_{0|t}}\mathfrak{L}(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y}),$
@@
(13) $\mathfrak{L}_{\text{Motion}}(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y})=\Sigma_{n}\Sigma_{l}m_{nl}\|\mathfrak{R}(g_{\bm{\psi}}(\bm{z}_{0|t}))_{nl}-y_{nl}\|_{2},$
@@
Discrete M2DM (Kong et al., 2023) $0.497^{\pm.003}$ $0.682^{\pm.002}$ $0.763^{\pm.003}$ $0.352^{\pm.005}$ $3.134^{\pm.010}$ $9.926^{\pm.073}$ $3.587^{\pm.072}$
AttT2M (Zhong et al., 2023) $0.499^{\pm.003}$ $0.690^{\pm.002}$ $0.786^{\pm.002}$ $0.112^{\pm.006}$ $3.038^{\pm.007}$ $9.700^{\pm.090}$ $2.452^{\pm.051}$
T2M-GPT (Zhang et al., 2023b) $0.492^{\pm.003}$ $0.679^{\pm.002}$ $0.775^{\pm.002}$ $0.141^{\pm.005}$ $3.121^{\pm.009}$ $9.722^{\pm.081}$ $1.831^{\pm.048}$
MoMask (Guo et al., 2024) $0.521^{\pm.002}$ $0.713^{\pm.002}$ $0.807^{\pm.002}$ $0.045^{\pm.002}$ $2.958^{\pm.008}$ $1.241^{\pm.040}$
DiverseMotion (Lou et al., 2023) $0.496^{\pm.004}$ $0.687^{\pm.004}$ $0.783^{\pm.003}$ $0.070^{\pm.004}$ $3.063^{\pm.011}$ $9.551^{\pm.068}$ $2.062^{\pm.079}$
MMM (Pinyoanuntapong et al., 2024b) $0.504^{\pm.003}$ $0.696^{\pm.003}$ $0.794^{\pm.002}$ $0.080^{\pm.003}$ $2.998^{\pm.007}$ $9.411^{\pm.058}$ $1.164^{\pm.041}$
ParCo (Zou et al., 2024) $0.515^{\pm.003}$ $0.706^{\pm.003}$ $0.801^{\pm.002}$ $0.109^{\pm.003}$ $2.927^{\pm.008}$ $9.576^{\pm.088}$ $1.382^{\pm.060}$
BAMM (Pinyoanuntapong et al., 2024a) $0.525^{\pm.002}$ $0.720^{\pm.003}$ $0.814^{\pm.003}$ $0.055^{\pm.002}$ $2.919^{\pm.008}$ $9.717^{\pm.089}$ $1.687^{\pm.051}$
Continuous (data space) MotionDiffuse (Zhang et al., 2024) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
MDM (Tevet et al., 2023) $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $2.799^{\pm.072}$
Fg-T2M (Wang et al., 2023) $0.492^{\pm.002}$ $0.683^{\pm.003}$ $0.783^{\pm.002}$ $0.243^{\pm.019}$ $3.109^{\pm.007}$ $9.278^{\pm.072}$ $1.614^{\pm.049}$
Continuous (latent) MLD (Chen et al., 2023) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $\bm{2.413^{\pm.079}}$
MotionLCM (Dai et al., 2024) $0.502^{\pm.003}$ $0.698^{\pm.002}$ $0.798^{\pm.002}$ $0.304^{\pm.012}$ $3.012^{\pm.007}$ $9.607^{\pm.066}$ $2.259^{\pm.092}$
@@
Table 1. Comparison with state-of-the-art methods on HumanML3D (Guo et al., 2022). Note that discrete representations do not allow for training-free motion editing; therefore, methods based on VQ-based latent representations (Discrete) are grayed out. The best scores for each metric in the methods using VAE-based latent representations (Continuous (latent)) are highlighted in bold.
4. Experiments

### v4

@@ 3.2.2 From GAN- to SAN-based discriminator
We apply the slicing adversarial network (SAN) framework [36] to further enhance the motion VAE, based on a prior report showing SAN-based models perform better than the GAN counterparts. The impact of this replacement on motion generation is discussed in Section 4.
3.2.3 Architectures
@@
We use a standard CNN-based architecture in the motion VAE encoder $q_{\bm{\eta}}$, decoder $g_{\bm{\psi}}$ and discriminator $f_{\bm{\psi}}$, consisting of 1D convolution, a residual block, and Leaky ReLU. For temporal downsampling and upsampling, we use stride 2 convolution and nearest interpolation, respectively. Specifically, the motion sequence $\bm{x}\in\mathbb{R}^{N\times L}$ is encoded into a latent vector $\bm{z}\in\mathbb{R}^{d_{z}\times d_{l}}$ with downsampling ratio of $d_{l}=L/4$. This architecture is inspired by [8, 44].
3.3 Stage 2: Motion Latent Diffusion
3.3.1 Text-conditional motion generation
@@
In this section, we train a text-conditioned diffusion model on the low-dimensional motion latent space obtained by the autoencoder learned in the stage 1 (Section 3.2). Using the trained model, we perform motion generation conditioned on text. First, we define a time-dependent sequence $\bm{z}_{0},\bm{z}_{1},\dots,\bm{z}_{t},\dots,\bm{z}_{T}\in\mathbb{R}^{d_{z}\times d_{l}}$ (starting from the VAE encoder output $\bm{z}_{0}=\bm{z}\sim q_{\bm{\eta}}(\bm{z}|\bm{x})$), which is derived from the following Markov diffusion process in the latent space:$q(\bm{z}_{t}|\bm{z}_{t-1})=\mathcal{N}(\sqrt{\alpha}_{t}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$ where $T>0$ and the constant $\alpha_{t}\in(0,1)$ is a pre-defined noise-scheduling parameter that determines the forward process. The forward process allows for the sampling of $\bm{z}_{t}$ at an arbitrary time step $t$ in a closed form: $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}\coloneq\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$.
@@
As our goal is text-to-motion generation, our interest is in the conditional distribution $p(\bm{z}|\bm{c})$ given the text prompt $\bm{c}$. Here, similar to many text-conditioned latent diffusion models [32, 3], we train the conditional model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ conditioned on the output of a text encoder $\tau(\bm{c})$, using the following objective function:
@@
$\mathcal{J}_{\text{cLDM}}(\theta)=\mathbb{E}_{\{\bm{z}_{0},\bm{c}\},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))\|_{2}^{2}\right],$ (8)
@@
where $\bm{z}_{0}$ and $\bm{c}$ are drawn from the joint empirical distribution. In addition, as done in prior works, we adopt classifier-free guidance [15] and train the model unconditionally, i.e., without a text prompt, with a certain probability during training.
@@
During inference, the trained diffusion model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ is used to generate $\bm{z}_{0}$ through a denoising process conditioned on the text prompt $\bm{c}$. We adopt the sampling scheme of DDIM [33] with trailing sample steps [22], in which each sampling step is defined as:
@@
$\ +\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))+\sigma_{t}\epsilon,$ (9)
@@
where $\sigma_{t}>0$ determines the stochasticity of the sampling process, and the sampling process becomes deterministic when $\sigma_{t}=0$. The part $(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c})))/\sqrt{\bar{\alpha}_{t}}$ in the first term corresponds to a direct estimate of the clean latent $\bm{z}_{0}$ from the noisy sample $\bm{z}_{t}$ using the diffusion model based on Tweedie’s formula [7]; this estimate is denoted as $\bm{z}_{0|t}$. In actual inference, we use the estimated $z_{0}$ and the VAE decoder trained in stage 1 to obtain $g_{\bm{\psi}}(\bm{z}_{0})$ as the generated motion sequences. Variable length motion generation is then achieved by clipping a part of $g_{\bm{\psi}}(\bm{z}_{0})$ where the activation variable $a^{i}$ introduced in Section 3.1 satisfies $a^{i}<\delta\ (0<\delta<1)$. The effect of this approach is discussed in Section 4.1.
3.3.2 Architecture
@@
We employ a diffusion transformer (DiT)-based architecture in our stage 2 model. The transformer used follows a standard structure of stacked blocks consisting of an attention layer and gated multilayer perceptrons (MLP) connected in series, with skip connections around each. Additionally, layer normalization is employed on the inputs of both the attention layer and the MLP. At the input and output of the transformer, linear mapping is used to convert from the latent dimension of the stage 1 model to the embedded dimension of the transformer. Text CLIP [31] embedding is also added as input to the transformer, along with an embedding describing the current time step of the diffusion process. This architecture is inspired by [9], which established SOTA performance in audio generation.
3.4 Controllable Motion Generation on Latent Diffusion Sampling
@@
In this section, we present a guided generation framework that leverages the pre-trained motion latent diffusion model (Section 3.3) for conditional motion generation and editing tasks without extra training. Major training-free methods (e.g, [4, 43, 14]) are based on the fact that the conditional score function can be decomposed into two additive terms: the unconditional score function and the log-likelihood term. Specifically, for a new condition $\bm{y}$, we have $\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t}|\bm{c},\bm{y})=\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t}|\bm{c})+\nabla_{\bm{z}_{t}}\log p(\bm{y}|\bm{z}_{t},\bm{c}),$ which is derived from Bayes’ rule. The conditional generation based on the aforementioned property can be regarded as a sequential procedure implemented as follows. First, a denoised sample $\bm{z}_{t-1}$ is obtained from $\bm{z}_{t}$ by the sampling step in Equation (9), without considering the given condition $\bm{y}$ (the first term in Equation (9)). Subsequently, $\bm{z}_{t-1}$ is further updated using the gradient of the log-likelihood term with respect to $\bm{z}_{t}$ (the second term).
@@
The challenge here is that the log-likelihood term is computed based on noisy samples $\bm{z}_{t}$. In the classifier-guided diffusion [35], time-dependent classifiers for $\bm{z}_{t}$ have to be trained, which requires additional training. In contrast, in training-free methods, this term is approximated with the current clean data estimate $\bm{z}_{0|t}$ and a loss function $\mathfrak{L}(\bm{x};\bm{y})$ defined for clean data. This loss function can be flexibly set depending on the task. For example, in inverse problems of the form $\bm{y}=\mathcal{A}(\bm{x})$, where $\mathcal{A}$ is a differentiable function with respect to $\bm{x}$, the loss function can be set as $\|\bm{y}-\mathcal{A}(\bm{x}_{0|t})\|_{2}^{2}$, where $\bm{x}_{0|t}=g_{\bm{\psi}}(\bm{z}_{0|t})$ is a clean data estimate in the original data domain and obtained through the VAE decoder. Here, we adopt MPGD [14], a fast yet high-quality guidance method applicable to latent diffusion models. Following the denoising step, it updates the denoised sample based on the loss function $\mathfrak{L}(\bm{x};\bm{y})$ as follows:
$\bm{z}_{t-1}\leftarrow\bm{z}_{t-1}-\rho_{t}\sqrt{\bar{\alpha}_{t-1}}\nabla_{\bm{z}_{0|t}}\mathfrak{L}(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y}),$ (10)
@@
$\mathfrak{L}_{\text{Motion}}(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y})=\Sigma_{n}\Sigma_{l}m_{nl}\|\mathfrak{R}(g_{\bm{\psi}}(\bm{z}_{0|t}))_{nl}-y_{nl}\|_{2},$ (11)
@@
Discrete M2DM [21] $0.497^{\pm.003}$ $0.682^{\pm.002}$ $0.763^{\pm.003}$ $0.352^{\pm.005}$ $3.134^{\pm.010}$ $9.926^{\pm.073}$ $3.587^{\pm.072}$
AttT2M [48] $0.499^{\pm.003}$ $0.690^{\pm.002}$ $0.786^{\pm.002}$ $0.112^{\pm.006}$ $3.038^{\pm.007}$ $9.700^{\pm.090}$ $2.452^{\pm.051}$
T2M-GPT [44] $0.492^{\pm.003}$ $0.679^{\pm.002}$ $0.775^{\pm.002}$ $0.141^{\pm.005}$ $3.121^{\pm.009}$ $9.722^{\pm.081}$ $1.831^{\pm.048}$
MoMask [13] $0.521^{\pm.002}$ $0.713^{\pm.002}$ $0.807^{\pm.002}$ $0.045^{\pm.002}$ $2.958^{\pm.008}$ $1.241^{\pm.040}$
DiverseMotion [24] $0.496^{\pm.004}$ $0.687^{\pm.004}$ $0.783^{\pm.003}$ $0.070^{\pm.004}$ $3.063^{\pm.011}$ $9.551^{\pm.068}$ $2.062^{\pm.079}$
MMM [29] $0.504^{\pm.003}$ $0.696^{\pm.003}$ $0.794^{\pm.002}$ $0.080^{\pm.003}$ $2.998^{\pm.007}$ $9.411^{\pm.058}$ $1.164^{\pm.041}$
ParCo [49] $0.515^{\pm.003}$ $0.706^{\pm.003}$ $0.801^{\pm.002}$ $0.109^{\pm.003}$ $2.927^{\pm.008}$ $9.576^{\pm.088}$ $1.382^{\pm.060}$
BAMM [28] $0.525^{\pm.002}$ $0.720^{\pm.003}$ $0.814^{\pm.003}$ $0.055^{\pm.002}$ $2.919^{\pm.008}$ $9.717^{\pm.089}$ $1.687^{\pm.051}$
Continuous (raw data) MotionDiffuse [46] $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
MDM [37] $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $2.799^{\pm.072}$
Fg-T2M [38] $0.492^{\pm.002}$ $0.683^{\pm.003}$ $0.783^{\pm.002}$ $0.243^{\pm.019}$ $3.109^{\pm.007}$ $9.278^{\pm.072}$ $1.614^{\pm.049}$
Continuous (latent) MLD [3] $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $\bm{2.413^{\pm.079}}$
MotionLCM [6] $0.502^{\pm.003}$ $0.698^{\pm.002}$ $0.798^{\pm.002}$ $0.304^{\pm.012}$ $3.012^{\pm.007}$ $9.607^{\pm.066}$ $2.259^{\pm.092}$
@@
Table 1: Comparison with state-of-the-art methods on HumanML3D dataset. Note that discrete representations do not allow for training-free motion editing; therefore, methods based on VQ-based latent representations (Discrete) are grayed out. The best scores for each metric in the methods using VAE-based latent representations (Continuous (latent)) are highlighted in bold.
4 Experiments